I need major help
solve each of the following quadiatic equations by completly the square.
x^2-6x-3=0
and 2x^2+10x+11=0
thanks a million

Answer by vertciel(183)

1. x^2 - 6x - 3 = 0
x^2 - 6x = 3
x^2 - 6x + 9 = 3 - 9
x^2 - 6x + 9 = -6
x^2 - 6x + 15 = 0
Use the quadratic equation to find the x values.

2. 2x^2 + 10x + 11 = 0
2x^2 + 10x = -11
2x^2 + 10x + 25 = - 11 - 25
2x^2 + 10x + 25 = -36
2x^2 + 10x + 61 = 0
Use the quadratic equation to find x.